Find the slope and intercept of the line 3x – 2y = -8.

Answer by Alan3354(69443)

Find the slope and intercept of the line 3x – 2y = -8.
-------
Put the equation into slope-intercept form, y = mx + b. That means solve for y.
3x – 2y = -8
2y = 3x + 8
y = (3/2)x + 4
------------
Slope, m, = 3/2
y-intercept, b, = 4